# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix F $-\ell$ Achievability Analysis for Near-Constant Weight Designs
Combining the bounds in (275) and (276), without accounting for the summation over $d$, a single summand in (269) corresponding to a specific $d$ can be bounded by
$\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\Big{[}\underbrace{e^{\nu}\cdot D(de^{-\nu}\|e^{-\nu})+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}}_{:=g(C,\zeta,d,\rho,\nu)}+o(1)\Big{]}\Big{)}.$ (277)
@@
Note that $g(C,\zeta,d,\rho,\nu)$ here has appeared in the proof of converse bound; see (219).
@@
$d^{*}=\operatorname*{arg\,min}_{d}\leavevmode\nobreak\ g(C,\zeta,d,\rho,\nu),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to }|C(2\zeta-1)|<d\leq e^{\nu},$ (279)
where we relax the range of $d$ from $\frac{C(2\zeta-1)}{1-2\rho}=d_{0}\leq d\leq e^{\nu}$ to $|C(2\zeta-1)|\leq d\leq e^{\nu}$ so that it matches the corresponding development in the converse proof (see (221)). We again write $A:=C(1-2\zeta)$, and note that $d^{*}$ in (279) has been pinpointed in (228). Therefore, the term in (269) is bounded by
$O\Big{(}\frac{\ell n}{k}\Big{)}\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}[g(C,\zeta,d^{*},\rho,\nu)+o(1)]\Big{)}=\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}[\underbrace{g(C,\zeta,d^{*},\rho,\nu)}_{:=f_{2}(C,\zeta,\rho,\nu)}+o(1)]\Big{)},$ (280)
where $f_{2}(C,\zeta,\rho,\nu)$ coincides with the one in the proof of near-constant weight design converse bound; see (231) in Appendix E.
Comparing the bounds in (274) and (280): Observe that (279) and $C(2\zeta-1)\leq d_{0}\leq e^{\nu}$ give
$f_{2}(C,\zeta,\rho,\nu)=g(C,\zeta,d^{*},\rho,\nu)\leq g(C,\zeta,d_{0},\rho,\nu)=e^{\nu}D(d_{0}e^{-\nu}\|e^{-\nu}),$ (281)
where the last equality follows from $D(\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\|\rho)=0$ when $d=d_{0}$. Thus, the bound in (280) dominates the one in (274), and by substituting into (266)–(269), we obtain
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}f_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}$ (282)
@@
Case 2: $d_{0}<0$ (i.e., $C(2\zeta-1)<0$). In this case, the term in (266) vanishes, and we focus on bounding the term in (269). As in Case 1, given $N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}$ we have $\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(\frac{d\nu e^{-\nu}\ell n}{k},\rho)$, and so we have the summand being 00 if $\frac{d+C(1-2\zeta)}{2}>d$, i.e., $d<C(1-2\zeta)$. Therefore, we can restrict the summation over $d$ in (269) to
$\mathcal{D}=\Big{\{}C(1-2\zeta)\leq d\leq e^{\nu}:\frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\Big{\}}.$ (283)
Then similarly to Case 1, we bound the $O(\frac{\ell n}{k})$ summands by their common upper bound as per (278) with $d^{*}$ given in (279), and obtain the same bound as (282).
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq 1,$ (284)
@@
Combining the Cases 1-3: We recall that $f_{2}(C,\zeta,\rho,\nu)$ is given in (280) and define
$\hat{f}_{2}(C,\zeta,\rho,\nu)=\begin{cases}f_{2}(C,\zeta,\rho,\nu)&\text{when }C(2\zeta-1)\leq 1-2\rho\\ 0&\text{when }C(2\zeta-1)>1-2\rho.\end{cases},$ (285)
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}.$ (286)
Additionally, we can verify that $f_{2}(C,\zeta,\rho,\nu)=0$ when $C(2\zeta-1)=1-2\rho$: Using (228) and some simple algebra, we find that $d^{*}=1$ holds when $A=C(1-2\zeta)=2\rho-1$; substituting this into $g(C,\zeta,d,\rho,\nu)$ in (277) yields
$g(C,\zeta,1,\rho,\nu)=D\Big{(}\frac{1+C(1-2\zeta)}{2}\big{\|}\rho\Big{)}=0.$ (287)
@@
((C2) holds for (,$\mathbb{P}\Big{(}\text{{(C2)} holds for }(\ell,C,\zeta)\Big{)}$ (288)
@@
$\quad\leq\binom{p}{\ell}\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}+k^{-10\ell}$ (290)
$\quad\leq\exp\Big{(}(1+o(1))\ell\log\frac{p}{\ell}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}+k^{-10\ell},$ (291)
where (289) holds because under the condition (261) we have that (264) is a necessary condition for “(C2) holds for $(\ell,C,\zeta)$”, and in (290) we apply a union bound over no more than $\binom{p}{\ell}$ possibilities of $\mathcal{J}^{\prime}$ to account for the maximum in (264), and recall that (261) holds with probability at least $1-k^{-10\ell}$. Therefore, for given $(\ell,C,\zeta)$ we have identified the following condition for (C2) to hold (where we now make the role of $\mathscr{A}_{1}$ explicit):
$\frac{\ell\nu ne^{-\nu}}{k}\hat{f}_{2}(C,\zeta,\rho,\nu)\geq(1+\eta_{2})\ell\log\frac{p}{\ell}\leavevmode\nobreak\ \text{ for some }\eta_{2}>0$ (292)
[…]
$\quad\quad\leq\exp\Big{(}-\frac{\eta_{2}}{2}\ell\log\frac{p}{\ell}\Big{)}+k^{-10\ell}\leq\hat{P}_{\ell}:=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k},\end{cases}$ (293)
since $\exp(-\frac{\eta^{2}}{2}\ell\log\frac{p}{\ell})$ is an upper bound on (291) when (292) holds.
@@
  • For given $\ell$ and any feasible $(C,\zeta)$, by (247)–(248), if
    $n\geq\frac{(1+\eta_{1})k\log\frac{k}{\ell}}{\nu e^{-\nu}f_{1}(C,\zeta,\rho,\nu)}$ (294)
    holds for some $\eta_{1}>0$ with $f_{1}(C,\zeta,\rho,\nu)$ being defined in (246), then (C1) holds for the given $(\ell,C,\zeta)$ with probability at most $\hat{P}_{\ell}$.
  • For given $\ell$ and any feasible $(C,\zeta)$, by (292)–(293), if
    $n\geq\frac{(1+\eta_{2})k\log\frac{p}{\ell}}{\nu e^{-\nu}[\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)]}$ (295)
    holds for some $\eta_{2}>0$ with $\hat{f}_{2}(C,\zeta,\rho,\nu)$ being given in (285) and (280), then the probability of “(C2) holds for the given $(\ell,C,\zeta)$ and $\mathscr{A}_{1}$ holds” is no more than $\hat{P}_{\ell}$.
@@
$n\geq(1+\eta)\frac{k}{\nu e^{-\nu}}\max_{\begin{subarray}{c}C\in(0,e^{\nu})\\ \zeta\in(0,1)\end{subarray}}\min\Big{\{}\frac{\log\frac{k}{\ell}}{f_{1}(C,\zeta,\rho,\nu)},\frac{\log\frac{p}{\ell}}{\hat{f}_{2}(C,\zeta,\rho,\nu)}\Big{\}}$ (298)
@@
($\mathbb{P}\big{(}\{\text{({C1}) and ({C2}) simultaneously hold for $(\ell,C,\zeta)\}$}\,\cap\,\mathscr{A}_{1}\big{)}\leq 2\hat{P}_{\ell}.$
Since it is unlikely for (C1) to hold when $C>e^{\nu}$, for the restricted MLE decoder in (30) we obtain that
@@
($\leq\mathbb{P}\Big{(}\{\textbf{(C1)}\text{ and }\textbf{(C2)}\leavevmode\nobreak\ \text{hold for some feasible }(C,\zeta)\in[0,e^{\nu}]\times[0,1]\}\cap\mathscr{A}_{1}\text{ holds}\Big{)}$ (299)
$\leq O((\ell\log k)^{2})\hat{P}_{\ell},$ (300)
where (299) holds due to Lemma 3(b), and (300) follows from a union bound over all feasible $(C,\zeta)$ under the given $\ell\in[1,\frac{k}{\log k}]$.
Bounding the overall failure probability for $\ell\in[1,\frac{k}{\log k}]$: We further take a union bound over all $\ell\in[1,\frac{k}{\log k}]$. We assume that the condition (298) holds for all $\ell\in[1,\frac{k}{\log k}]$; by the same reasoning as (157)–(158), this assumption amounts to
$n\geq\frac{(1+\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),\hat{f}_{2}(C,\zeta,\rho,\nu)\}}$ (301)
for some $\eta>0$. Then, if (301) holds, by union bound we can bound the overall failure probability as
$\mathbb{P}({\rm err})$ $\leq\mathbb{P}(\mathscr{A}_{1}^{c})+\sum_{\ell=1}^{k/\log k}\mathbb{P}\Big{(}\Big{\{}|s\setminus\widehat{S}^{\prime}|=\ell\Big{\}}\cap\mathscr{A}_{1}\Big{)}$ (302)
$\leq o(1)+\sum_{\ell=1}^{k/\log k}O((\ell\log k)^{2})\hat{P}_{\ell}$ (303)
$\leq o(1)+\Big{[}\sum_{\ell=1}^{\log k}(\ell\log k)^{-3}+\sum_{\ell=\log k}^{k/\log k}k^{-3}\Big{]}=o(1),$ (304)
where (302) follows from a union bound, and we use (300) in (303) and then substitute $\hat{P}_{\ell}$ (given by (293)) in (304). In conclusion, if (301) holds for some $\eta>0$, then restricted MLE has $o(1)$ probability of failure.
Simplifying $\hat{f}_{2}(C,\zeta,\rho,\nu)$ to $f_{2}(C,\zeta,\rho,\nu)$: Recall that $\hat{f}_{2}(C,\zeta,\rho,\nu)$ is given in (285). In this step, we show that the minimum over $(C,\zeta)$ in (301) is not attained in the domain of $C(2\zeta-1)>1-2\rho$, thus we can safely simplify $\hat{f}_{2}(C,\zeta,\rho,\nu)$ to $f_{2}(C,\zeta,\rho,\nu)$ in the threshold (301). To this end, note that $C(2\zeta-1)>1-2\rho$ is equivalent to
@@
and we note that $\hat{f}_{2}(C,\zeta,\rho,\nu)=0$ holds for any $\zeta\geq\zeta^{\prime}$ (see (285) and recall that $\hat{f}_{2}(C,\zeta,\rho,\nu)$ is continuous), and that $f_{1}(C,\zeta,\rho,\nu)$ is monotonically increasing with respect to $\zeta$ in $[\frac{1}{2},1)$ (see (246)). Therefore, for any given $(C,\zeta)$ with $C(2\zeta-1)>1-2\rho$ we can also consider $(C,\zeta^{\prime})$ that falls in the case of $C(2\zeta-1)\leq 1-2\rho$, and we can compare the values of $f_{1}(C,\zeta,\rho,\nu)$ and $\hat{f}_{2}(C,\zeta,\rho,\nu)$ as
@@
$\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),\hat{f}_{2}(C,\zeta,\rho,\nu)\Big{\}}\geq\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta^{\prime},\rho,\nu),\hat{f}_{2}(C,\zeta^{\prime},\rho,\nu)\Big{\}}.$ (305)
Thus, the minimum over $(C,\zeta)$ in (301) is attained at the case $C(2\zeta-1)\leq 1-2\rho$, in which $\hat{f}_{2}(C,\zeta,\rho,\nu)=f_{2}(C,\zeta,\rho,\nu)$. Then similarly to (165)–(174) we have
$\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),\hat{f}_{2}(C,\zeta,\rho,\nu)\Big{\}}$ (306)
$\geq$ $\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),f_{2}(C,\zeta,\rho,\nu)\Big{\}},$ (307)
implying that to ensure (301) it suffices to have
$n\geq\frac{(1+\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),f_{2}(C,\zeta,\rho,\nu)\}}.$ (308)
This establishes the second term in (11).
Appendix G Technical Lemmas for Near-Constant Weight Designs
Compared to the Bernoulli design, an additional challenge in the near-constant weight design is that a test may receive multiple placements from a single item. We present two high-probability events that prove useful in overcoming this difficulty.
Lemma 5.
([16, Prop. II.3] and [16, Eq. (40)]) Under the near-constant design where each item is uniformly placed (with replacement) to $\Delta=\frac{\nu n}{k}$ tests, we have the following:

### 23 Oct 2024 v2

@@ Appendix F $-\ell$ Achievability Analysis for Near-Constant Weight Designs
Combining the bounds in (F.33) and (F.34), without accounting for the summation over $d$, a single summand in (F.27) corresponding to a specific $d$ can be bounded by
$\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\Big{[}\underbrace{e^{\nu}\cdot D(de^{-\nu}\|e^{-\nu})+d\cdot D\Big{(}\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\big{\|}\rho\Big{)}}_{:=g(C,\zeta,d,\rho,\nu)}+o(1)\Big{]}\Big{)}.$ (F.35)
@@
Note that $g(C,\zeta,d,\rho,\nu)$ here has appeared in the proof of converse bound; see (E.44).
@@
$d^{*}=\operatorname*{arg\,min}_{d}\leavevmode\nobreak\ g(C,\zeta,d,\rho,\nu),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{subject to }|C(2\zeta-1)|<d\leq e^{\nu},$ (F.37)
and where we relax the range of $d$ from $\frac{C(2\zeta-1)}{1-2\rho}=d_{0}\leq d\leq e^{\nu}$ to $|C(2\zeta-1)|\leq d\leq e^{\nu}$ so that it matches the corresponding development in the converse proof (see (E.46)). We again write $A:=C(1-2\zeta)$, and note that $d^{*}$ in (F.37) has been pinpointed in (E.53). Therefore, the term in (F.27) is bounded by
$O\Big{(}\frac{\ell n}{k}\Big{)}\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}[g(C,\zeta,d^{*},\rho,\nu)+o(1)]\Big{)}=\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}[\underbrace{g(C,\zeta,d^{*},\rho,\nu)}_{:=f_{2}(C,\zeta,\rho,\nu)}+o(1)]\Big{)},$ (F.38)
where $f_{2}(C,\zeta,\rho,\nu)$ coincides with the one in the proof of near-constant weight design converse bound; see (E.56) in Appendix E.
Comparing the bounds in (F.32) and (F.38): Observe that (F.37) and $C(2\zeta-1)\leq d_{0}\leq e^{\nu}$ give
$f_{2}(C,\zeta,\rho,\nu)=g(C,\zeta,d^{*},\rho,\nu)\leq g(C,\zeta,d_{0},\rho,\nu)=e^{\nu}D(d_{0}e^{-\nu}\|e^{-\nu}),$ (F.39)
where the last equality follows from $D(\frac{1}{2}+\frac{C(1-2\zeta)}{2d}\|\rho)=0$ when $d=d_{0}$. Thus, the bound in (F.38) dominates the one in (F.32), and by substituting into (F.26)–(F.27), we obtain
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}f_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}$ (F.40)
@@
Case 2: $d_{0}<0$ (i.e., $C(2\zeta-1)<0$). In this case, the term in (F.26) vanishes, and we focus on bounding the term in (F.27). As in Case 1, given $N_{0,\mathcal{J}^{\prime}}=\frac{d\nu e^{-\nu}\ell n}{k}$ we have $\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(\frac{d\nu e^{-\nu}\ell n}{k},\rho)$, and so we have the summand being 00 if $\frac{d+C(1-2\zeta)}{2}>d$, i.e., $d<C(1-2\zeta)$. Therefore, we can restrict the summation over $d$ in (F.27) to
$\mathcal{D}=\Big{\{}C(1-2\zeta)\leq d\leq e^{\nu}:\frac{d\nu e^{-\nu}\ell n}{k}\in\mathbb{Z}\Big{\}}.$ (F.41)
Then similarly to Case 1, we bound the $O(\frac{\ell n}{k})$ summands by their common upper bound as per (F.36) with $d^{*}$ given in (F.37), and obtain the same bound as (F.40).
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq 1,$ (F.42)
@@
Combining the Cases 1-3: We recall that $f_{2}(C,\zeta,\rho,\nu)$ is given in (F.38) and define
$\hat{f}_{2}(C,\zeta,\rho,\nu)=\begin{cases}f_{2}(C,\zeta,\rho,\nu)&\text{when }C(2\zeta-1)\leq 1-2\rho\\ 0&\text{when }C(2\zeta-1)>1-2\rho.\end{cases},$ (F.43)
@@
$\mathbb{P}\Big{(}\widetilde{G}_{\mathcal{J}^{\prime},1}-\frac{1}{2}N_{0,\mathcal{J}^{\prime}}\geq\big{(}\frac{1}{2}-\zeta-o(1)\big{)}\frac{Cn\nu e^{-\nu}\ell}{k}\Big{)}\leq\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}.$ (F.44)
Additionally, we can verify that $f_{2}(C,\zeta,\rho,\nu)=0$ when $C(2\zeta-1)=1-2\rho$: Using (E.53) and some simple algebra, we find that $d^{*}=1$ holds when $A=C(1-2\zeta)=2\rho-1$; substituting this into $g(C,\zeta,d,\rho,\nu)$ in (F.35) yields
$g(C,\zeta,1,\rho,\nu)=D\Big{(}\frac{1+C(1-2\zeta)}{2}\big{\|}\rho\Big{)}=0.$ (F.45)
@@
((C2) holds for (,$\mathbb{P}\Big{(}\text{{(C2)} holds for }(\ell,C,\zeta)\Big{)}$ (F.46)
@@
$\quad\leq\binom{p}{\ell}\exp\Big{(}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}+k^{-10\ell}$ (F.48)
$\quad\leq\exp\Big{(}(1+o(1))\ell\log\frac{p}{\ell}-\frac{\ell\nu ne^{-\nu}}{k}\big{[}\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)\big{]}\Big{)}+k^{-10\ell},$ (F.49)
where (F.47) holds because under the condition (F.23) we have that (F.24) is a necessary condition for “(C2) holds for $(\ell,C,\zeta)$”, and in (F.48) we apply a union bound over no more than $\binom{p}{\ell}$ possibilities of $\mathcal{J}^{\prime}$ to account for the maximum in (F.24), and recall that (F.23) holds with probability at least $1-k^{-10\ell}$. Combining these developments, we deduce that the following implication holds for any $\eta_{2}>0$ and sufficiently large $p$ (where we now make the role of $\mathscr{A}_{1}$ explicit):
$\frac{\ell\nu ne^{-\nu}}{k}\hat{f}_{2}(C,\zeta,\rho,\nu)\geq(1+\eta_{2})\ell\log\frac{p}{\ell}$ (F.50)
($\Longrightarrow\leavevmode\nobreak\ \mathbb{P}\Big{(}\big{\{}\textbf{(C2)}\text{ holds for }(\ell,C,\zeta)\big{\}}\cap\mathscr{A}_{1}\Big{)}$ $\leq\exp\Big{(}-\frac{\eta_{2}}{2}\ell\log\frac{p}{\ell}\Big{)}+k^{-10\ell}$ (F.51)
$\leq\hat{P}_{\ell}:=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k},\end{cases}$ (F.52)
since $\exp(-\frac{\eta^{2}}{2}\ell\log\frac{p}{\ell})$ is an upper bound on (F.49) when (F.50) holds.
@@
  • For given $\ell$ and any feasible $(C,\zeta)$, by (F.13)–(F.14), if
    $n\geq\frac{(1+\eta_{1})k\log\frac{k}{\ell}}{\nu e^{-\nu}f_{1}(C,\zeta,\rho,\nu)}$ (F.53)
    holds for some $\eta_{1}>0$ with $f_{1}(C,\zeta,\rho,\nu)$ being defined in (F.12), then (C1) holds for the given $(\ell,C,\zeta)$ with probability at most $\hat{P}_{\ell}$.
  • For given $\ell$ and any feasible $(C,\zeta)$, by (F.50)–(F.52), if
    $n\geq\frac{(1+\eta_{2})k\log\frac{p}{\ell}}{\nu e^{-\nu}[\hat{f}_{2}(C,\zeta,\rho,\nu)+o(1)]}$ (F.54)
    holds for some $\eta_{2}>0$ with $\hat{f}_{2}(C,\zeta,\rho,\nu)$ being given in (F.43) and (F.38), then the probability of “(C2) holds for the given $(\ell,C,\zeta)$ and $\mathscr{A}_{1}$ holds” is no higher than $\hat{P}_{\ell}$.
@@
$n\geq(1+\eta)\frac{k}{\nu e^{-\nu}}\max_{\begin{subarray}{c}C\in(0,e^{\nu})\\ \zeta\in(0,1)\end{subarray}}\min\Big{\{}\frac{\log\frac{k}{\ell}}{f_{1}(C,\zeta,\rho,\nu)},\frac{\log\frac{p}{\ell}}{\hat{f}_{2}(C,\zeta,\rho,\nu)}\Big{\}}$ (F.55)
@@
($\mathbb{P}\big{(}\{\text{({C1}) and ({C2}) simultaneously hold for $(\ell,C,\zeta)\}$}\,\cap\,\mathscr{A}_{1}\big{)}\leq\hat{P}_{\ell}.$
Since it is impossible for (C1) to hold when $C>e^{\nu}$, for the restricted MLE decoder in (B.2) we obtain that
@@
($\leq\mathbb{P}\Big{(}\{\textbf{(C1)}\text{ and }\textbf{(C2)}\leavevmode\nobreak\ \text{hold for some feasible }(C,\zeta)\in[0,e^{\nu}]\times[0,1]\}\cap\mathscr{A}_{1}\text{ holds}\Big{)}$ (F.56)
$\leq O((\ell\log k)^{2})\hat{P}_{\ell},$ (F.57)
where (F.56) holds due to Lemma B.1(b), and (F.57) follows from a union bound over all feasible $(C,\zeta)$ under the given $\ell\in[1,\frac{k}{\log k}]$.
Bounding the overall failure probability for $\ell\in[1,\frac{k}{\log k}]$: We further take a union bound over all $\ell\in[1,\frac{k}{\log k}]$. Suppose that the condition (F.55) holds for all $\ell\in[1,\frac{k}{\log k}]$; by the same reasoning as (D.70)–(D.71), this amounts to
$n\geq\frac{(1+\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),\hat{f}_{2}(C,\zeta,\rho,\nu)\}}$ (F.58)
for some $\eta>0$. Then, if (F.58) holds, we can bound the overall failure probability as
$\mathbb{P}({\rm err})$ $\leq\mathbb{P}(\mathscr{A}_{1}^{c})+\sum_{\ell=1}^{k/\log k}\mathbb{P}\Big{(}\Big{\{}|s\setminus\widehat{S}^{\prime}|=\ell\Big{\}}\cap\mathscr{A}_{1}\Big{)}$ (F.59)
$\leq o(1)+\sum_{\ell=1}^{k/\log k}O((\ell\log k)^{2})\hat{P}_{\ell}$ (F.60)
$\leq o(1)+\Big{[}\sum_{\ell=1}^{\log k}(\ell\log k)^{-3}+\sum_{\ell=\log k}^{k/\log k}k^{-3}\Big{]}=o(1),$ (F.61)
where (F.59) follows from the union bound, (F.60) follows from (F.57), and in (F.61) we substitute $\hat{P}_{\ell}$ (given by (F.52)). In conclusion, if (F.58) holds for some $\eta>0$, then restricted MLE has $o(1)$ probability of failure.
Simplifying $\hat{f}_{2}(C,\zeta,\rho,\nu)$ to $f_{2}(C,\zeta,\rho,\nu)$: Recall that $\hat{f}_{2}(C,\zeta,\rho,\nu)$ is given in (F.43). In this step, we show that the minimum over $(C,\zeta)$ in (F.58) is not attained in the domain of $C(2\zeta-1)>1-2\rho$, thus we can safely simplify $\hat{f}_{2}(C,\zeta,\rho,\nu)$ to $f_{2}(C,\zeta,\rho,\nu)$ in the threshold (F.58). To this end, note that $C(2\zeta-1)>1-2\rho$ is equivalent to
@@
and we note that $\hat{f}_{2}(C,\zeta,\rho,\nu)=0$ holds for any $\zeta\geq\zeta^{\prime}$ (see (F.43) and recall that $\hat{f}_{2}(C,\zeta,\rho,\nu)$ is continuous), and that $f_{1}(C,\zeta,\rho,\nu)$ is monotonically increasing with respect to $\zeta$ in $[\frac{1}{2},1)$ (see (F.12)). Therefore, for any given $(C,\zeta)$ with $C(2\zeta-1)>1-2\rho$ we can also consider $(C,\zeta^{\prime})$ that falls in the case of $C(2\zeta-1)\leq 1-2\rho$, and we can compare the values of $f_{1}(C,\zeta,\rho,\nu)$ and $\hat{f}_{2}(C,\zeta,\rho,\nu)$ as
@@
$\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),\hat{f}_{2}(C,\zeta,\rho,\nu)\Big{\}}\geq\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta^{\prime},\rho,\nu),\hat{f}_{2}(C,\zeta^{\prime},\rho,\nu)\Big{\}}.$ (F.62)
Thus, the minimum over $(C,\zeta)$ in (F.58) is attained at the case $C(2\zeta-1)\leq 1-2\rho$, in which $\hat{f}_{2}(C,\zeta,\rho,\nu)=f_{2}(C,\zeta,\rho,\nu)$. Then similarly to (D.78)–(D.81) we have
$\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),\hat{f}_{2}(C,\zeta,\rho,\nu)\Big{\}}$ (F.63)
$\geq$ $\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\Big{\{}\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),f_{2}(C,\zeta,\rho,\nu)\Big{\}},$ (F.64)
implying that to ensure (F.58) it suffices to have
$n\geq\frac{(1+\eta)k\log\frac{p}{k}}{(1-\theta)\nu e^{-\nu}}\frac{1}{\min_{C\in(0,e^{\nu}),\zeta\in(0,1)}\max\{\frac{1}{\theta}f_{1}(C,\zeta,\rho,\nu),f_{2}(C,\zeta,\rho,\nu)\}}.$ (F.65)
This establishes the second term in (1.11).
Appendix G Technical Lemmas for Near-Constant Weight Designs ($-\ell$)
Compared to the Bernoulli design, an additional challenge in the near-constant weight design is that a test may receive multiple placements from a single item. We present two high-probability events that are useful in overcoming this difficulty.
Lemma G.1.
([15, Prop. II.3] and [15, Eq. (40)]) Under the near-constant design where each item is uniformly placed (with replacement) to $\Delta=\frac{\nu n}{k}$ tests, we have the following: